Among 15 clocks there are two defectives. In how many ways can an inspector chose


three of the clocks for inspection so that:


a. There is no restriction.


b. None of the defective clock is included.


c. Only one of the defective clocks is included.


d. Two of the defective clock is included.


1
Expert's answer
2022-04-11T14:59:45-0400

a."C^3_{15}=\frac{15!}{12!3!}=455"

b."C^3_{15-2}=\frac{13!}{10!3!}=286"

c. "C^1_2C^2_{13}=\frac{2!}{1!1!}\frac{13!}{11!2!}=156"

d."C^2_2C^1_{13}=1\cdot\frac{13!}{12!1!}=13"